OpenStudy (anonymous):

simplify 6/x+3 + x/x+3

OpenStudy (anonymous):

\[\frac{ 6 }{ x+3 }+\frac{ x }{ x+3 }\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 6+x }{ x+3 }\]